OpenStudy (anonymous):

How do I re-write this in standard form? : 1/2x + 1/2y = 6

OpenStudy (anonymous):

is this \[\frac{1}{2}x+\frac{1}{2}y=6\]

hero (hero):

yup

hero (hero):

I prefer to write it as: \[\frac{x}{2} + \frac{y}{2} = 6 = \frac{x+y}{2}\]

OpenStudy (anonymous):

x+y=12

hero (hero):

:)

OpenStudy (anonymous):

Oh thanks. Soo it's already simplified?

OpenStudy (anonymous):

i would write x+y=12 myself

OpenStudy (anonymous):

Okay, thank you guyss. :))